# Rethinking Multi-modal Object Detection from the Perspective of Mono-Modality Feature Learning

### v1

@@ Abstract
Multi-Modal Object Detection (MMOD), due to its stronger adaptability to various complex environments, has been widely applied in various applications. Extensive research is dedicated to the RGB-IR object detection, primarily focusing on how to integrate complementary features from RGB-IR modalities. However, they neglect the mono-modality insufficient learning problem that the decreased feature extraction capability in multi-modal joint learning. This leads to an unreasonable but prevalent phenomenon—Fusion Degradation, which hinders the performance improvement of the MMOD model. Motivated by this, in this paper, we introduce linear probing evaluation to the multi-modal detectors and rethink the multi-modal object detection task from the mono-modality learning perspective. Therefore, we construct an novel framework called M2D-LIF, which consists of the Mono-Modality Distillation (M2D) method and the Local Illumination-aware Fusion (LIF) module. The M2D-LIF framework facilitates the sufficient learning of mono-modality during multi-modal joint training and explores a lightweight yet effective feature fusion manner to achieve superior object detection performance. Extensive experiments conducted on three MMOD datasets demonstrate that our M2D-LIF effectively mitigates the Fusion Degradation phenomenon and outperforms the previous SOTA detectors.Corresponding author: Maoxun Yuan and Xingxing Wei. represent the equal contribution to this work.
@@ 1 Introduction
Recently, multi-modal object detection (MMOD) technology has been widely used in various safety-critical applications such as around-the-clock pedestrian detection for urban surveillance [10], object detection for autonomous driving [21] etc. Since relying solely on visible images will render objects invisible under limited illumination, utilizing both visible (RGB) images and infrared (IR) images has been widely used as one of the effective solutions to achieve full-time object detection tasks. Specifically, RGB modality can provide the texture of objects in daylight, while infrared modality can provide the outline of objects under poor lighting conditions. By exploring the complementary information between the two modalities, RGB-IR object detection [15, 40] can aggregate the advantages of each modality to achieve robust visual perception.
@@
To design advanced fusion strategies for MMOD task, previous works [17, 13, 29] reveal that the feature-level fusion methods outperforms both image-level and decision-level fusion methods. Building upon this finding, recent RGB-IR object detection methods mainly focus on designing complex feature fusion structures to address various challenges. For example, CSSA [2] and ICAFusion [25] introduce different attention mechanisms to better explore multi-modal complementarity. CALNet [7] proposes the selected cross-modal fusion module to address semantic conflict issue between different modalities. AR-CNN [41, 42] and TSFADet [34, 36] focus on solving the misalignment problem and attempt to improve object detection performance by aligning RGB and IR modality features. Although these methods have achieved encouraging object detection performance, they primarily focus on integrating complementary features from RGB-IR modalities while neglecting the mono-modality insufficient learning problem. This results in an unreasonable phenomenon that some objects can be detected by the mono-modal detector but missed by the corresponding multi-modal detector. We name this phenomenon as Fusion Degradation as illustrated in Figure 1. Besides, we statistics this phenomenon on three commonly used MMOD datasets and find that it is prevalent in different types of multi-modal detectors. Motivated by this, we rethink the multi-modal object detection task from the mono-modality learning perspective.
@@
In this paper, we construct M2D-LIF, an end-to-end framework for multi-modal object detection, to address the above problem. Different from the existing MMOD methods that design complex feature fusion modules, M2D-LIF facilitates the sufficient learning of mono-modality during multi-modal joint training and explore a lightweight yet effective feature fusion manner to achieve superior object detection performance. Inspired by knowledge distillation [8], we propose an Mono-Modality Distillation (M2D) method, which introduces the additional encoder pretrained on the mono-modality to distill the multi-modal encoders. To bridge the capability gap between the multi- and mono-modalities, we design the inner-modality and cross-modality distillation loss to jointly optimize the M2D-LIF during training. In this way, we can ensure the sufficient learning of mono-modal encoder in the multi-modality joint training. Besides, according to our observations in Section 3.1, we propose a Local Illumination-aware Fusion (LIF) module, which can dynamically set different weights for different illumination regions. The module ensures an explicitly complementary fusion to cooperate with M2D method, thereby enhancing both the accuracy and efficacy of the multi-modal object detection. Figure 1 shows that our framework can effectively solve the above problem. Our contributions in this paper are highlighted as follows:
@@
  • We introduce linear probing evaluation to the multi-modal detectors and identify the insufficient learning of mono-modality during training. To the best of our knowledge, it is the first time to rethink multi-modal object detection from mono-modality learning perspective.
@@
  • We present M2D-LIF, a pioneering method to improve mono-modality learning capabilities during multi-modal joint training, innovatively exploring a lightweight yet effective feature fusion framework for MMOD task.
@@
  • Extensive experiments on the three MMOD datasets demonstrate that our M2D-LIF outperforms the previous stat-of-the-art detectors and can be used as a effective feature fusion method in MMOD task.
@@ 2.1 RGB-Infrared Object Detection
However, above methods mainly focus on designing complex and parameter-intensive modules to achieve the so-called complementary fusion, while ignoring the insufficient representation of each modality caused by the multi-modal learning. Different from these approaches, in this paper, we conduct the RGB-IR object detection task from mono-modality feature learning perspective and propose a novel M2D-LIF framework that only introduces a lightweight feature fusion module to achieve superior object detection performance.
@@ 2.2 Knowledge Distillation for Object Detection
Knowledge Distillation (KD) [8] can transfer knowledge from a high-capacity teacher model to a compact student model, thereby reducing the complexity of the model while ensuring model performance. For object detection task, FRS [45] and PGD [31] leveraged foreground masks to enhance the ability of student model to capture critical object features. CWD [26] normalized activation maps of each channel to obtain soft probability maps and minimizing the KL divergence between these maps, enabling the student to focus on the most salient regions of each channel. Besides, PKD [1] imitated features with pearson correlation coefficient to focus on the relational information from the teacher and relax constraints on the magnitude of the features. Recently, CrossKD [28] introduced cross-head knowledge distillation by transferring the intermediate features from the student detection head to the teacher providing a diversified perspective. As for multi-modal object detection task, KD is usually used to transfer the knowledge from multi-modality features to mono-modality features. For example, distillation from Radar-Lidar to Radar [30], BEV-Lidar to BEV [6], and RGB-Infrared to RGB [20]. In our paper, unlike the above methods, we design a novel knowledge distillation method called Mono-Modality Distillation (M2D) to improve the feature representation of each backbone network in multi-modality join training.
Figure 2: Liner probing evaluation on the FLIR dataset. Three types of feature fusion methods are selected for comparison, such as naive addition (Halfway Fusion [18]), weighted-based (IWM [9]) and attention-based (CMX [39]).
@@ 3.1 Linear Probing Evaluation
In current MMOD framework, each modality is encoded by its corresponding backbone network and then a fusion module is utilized to obtain the fused features for downstream perception tasks. Although achieving the superior performance, we claim that this way will lead to the insufficient learning of each modality and thus fail to achieve optimal detection performance. To validate this point, we employ the linear probing evaluation. Specifically, we first train the mono-modal object detectors and the multi-modal object detectors respectively. For multi-modal object detectors, we select three popular feature fusion methods for comparison, such as Halfway Fusion [18] (naive addition), IWM [9] (weighted-based) and CMX [39] (attention-based). Note that, all detectors utilize the CSPDarknet53 [11] as the backbone network. After that, each backbone from both the mono- and multi-modality detectors is frozen and connected with the new detection head for training and testing. As shown in Figure 2, the evaluation results reveal that:
(1) The backbone networks from multi-modality join training are insufficient learning. After linear probing evaluation, we can observe that all backbone networks from multi-modality join training are worse than those from mono-modality training. This indicates that due to the existence of the fusion module, the learning ability of each backbone network is limited during multi-modal training.
@@
Figure 3: Overview of our $\rm M^{2}$-MFD framework. The proposed M2D is for enhancing mono-modal feature extraction capability of the backbone in multi-modal object detection task. The proposed LIF method can evaluate the mono-modal quality based on RGB image’s local illumination conditions. Noted that since our framework is designed from the mono-modality perspective, the design of the infrared branch and the RGB branch is symmetrical and has the same network structure. Therefore, only the image processing flow of the RGB modality is emphasized in the figure, and the infrared branch is omitted.
@@ 3.2 M2D-LIF Framework
From Section 3.1 we have observed that the recent MMOD methods still suffer from the insufficient learning of mono-modality features. To solve this issue, referring to the knowledge distillation technology, we consider utilizing the additional encoder pretrained on the mono-modality as the teacher model to distill the multi-modality backbone network during training. Thus, under the premise that the multi-modal encoder learning sufficiently, we design a novel weighted-based feature fusion method to further improve the performance. The overall framework (M2D-LIF) is illustrated in Figure 3, which mainly consists of ➊ Mono-Modality Distillation and ➋ Local Illumination-aware Fusion. Note that the IR branch is omitted in Figure 3 since it has the same network structure as the RGB branch.
Mono-Modality Distillation (M2D). The M2D aims to enhance the feature extraction capability of the multi-modal encoders, laying solid foundation for improving object detection performance. To bridge the capability gap between the multi- and mono-modal encoders, we first employ a pretrained teacher model of the same modality to distill the multi-modal backbone network using the inner-modality distillation loss $\mathcal{L}_{\text{IM}}$. To further enhance object-relevant feature extraction, we propose cross-modality distillation loss $\mathcal{L}_{\text{CM}}$, which leverages cross-modality salient object location priors to guide feature distillation.
@@
For the loss function $\mathcal{L}_{\text{IM}}$ represents the inner-modality feature distillation loss, which guides the backbone of the multi-modal object detection model to learn intermediate layer responses from a teacher backbone of the same modality. By minimizing $\mathcal{L}_{\text{IM}}$, the multi-modal backbone is encouraged to align with the feature responses of the teacher model. The loss is defined as follows:
@@
As for the loss function $\mathcal{L}_{\text{CM}}$ represents the cross-modality feature distillation loss, which integrates cross-modality target location priors. Specifically, in MMOD tasks, different modalities generally contain information about the same objects. We first employ an attention mechanism to extract the salient object feature attention map, which serves as the location prior. The attention map is then used as a mask to guide object-relevant feature learning. To maintain a lightweight design, we adopt the parameter-free attention method SimAM [32]. The attention map $\widetilde{\mathcal{M}}$ can be calculated as:
@@
where $B$ donates the predicted brightness map. We transform the RGB image to the $LAB$ color space and extract $L$ channel as the ground-truth $\widetilde{B}$ to supervise the brightness prediction, which is formulated as follows:
@@
where $W_{V}$ and $W_{I}$ represent the weight of the RGB and infrared modalities, respectively. The hyperparameter $\alpha$ is the threshold that determines the importance of RGB features and $\beta$ is the amplitude of the $W_{V}$. Thus, the boundary of $W_{V}$ and $W_{I}$ are:
@@
Finally, the LIF module uses the weight map $W_{V}$ and $W_{I}$ to perform element-wise weighted fusion for the multi-modal features. Therefore, the final fused feature can be represented as:
@@
where $\lambda_{M^{2}D}$ and $\lambda_{LI}$ are the hyper-parameters that control the balance between each loss.
@@
Table 1: Comparison of detection performance (mAP, in%) and computational cost (Params , FLOPs) between the LIF and other SOTA fusion methods on the FLIR (F), DroneVehicle (D) and LLVIP (L) datasets. The symbol indicates training with our M2D method. The Best results are highlighted by bold.
@@ 3.3 Why Using LIF Module for Feature Fusion
(1) robustness-and-reasonable. As analyzed in Section 3.1, the weighted-based method can serve as a competitive fusion way to improve performance. As a weighted-based fusion module, LIF is more robust to the insufficient learning of mono-modality encoders, thus achieving superior performance. Furthermore, unlike current weighted-based methods, LIF module provides reasonably fine-grained quality assessment for the mono-modality images and explicitly achieve the complementary fusion.
(2) lightweight-yet-effective. As shown in Table 1, although attention-based methods can adaptively calculate the required features between modalities, they typically rely on high computational complexity. Compare with other methods, our LIF only introduces an additional 0.1M parameters and 6.4 GFLOPs in computational cost to achieve the superior performance, which is more lightweight and effective.
Table 2: Ablation on our M2D-LIF framework with performance measured by mAP (%). The best results are highlighted in bold.
M2D LIF FLIR DroneVehicle LLVIP
43.5 66.9 68.2
45.1 69.2 70.7
46.1 70.6 70.8
Table 3: Ablation on different way of the weight map generation in the LIF module with performance measured by mAP (%). The best results are highlighted in bold.
Method FLIR DroneVehicle LLVIP
M2D-LIF (L-channel) 44.2 67.4 70.2
M2D-LIF (Ours) 46.1 70.6 70.8

### v2

@@ Abstract
Multi-Modal Object Detection (MMOD), due to its stronger adaptability to various complex environments, has been widely applied in various applications. Extensive research is dedicated to the RGB-IR object detection, primarily focusing on how to integrate complementary features from RGB-IR modalities. However, they neglect the mono-modality insufficient learning problem, which arises from decreased feature extraction capability in multi-modal joint learning. This leads to a prevalent but unreasonable phenomenon—Fusion Degradation, which hinders the performance improvement of the MMOD model. Motivated by this, in this paper, we introduce linear probing evaluation to the multi-modal detectors and rethink the multi-modal object detection task from the mono-modality learning perspective. Therefore, we construct a novel framework called M2D-LIF, which consists of the Mono-Modality Distillation (M2D) method and the Local Illumination-aware Fusion (LIF) module. The M2D-LIF framework facilitates the sufficient learning of mono-modality during multi-modal joint training and explores a lightweight yet effective feature fusion manner to achieve superior object detection performance. Extensive experiments conducted on three MMOD datasets demonstrate that our M2D-LIF effectively mitigates the Fusion Degradation phenomenon and outperforms the previous SOTA detectors. The codes are available at https://github.com/Zhao-Tian-yi/M2D-LIF.
footnotetext: *: Equal contribution. $\dagger$: Corresponding Author.
@@ 1 Introduction
Recently, multi-modal object detection (MMOD) technology has been widely used in various safety-critical applications such as around-the-clock pedestrian detection for urban surveillance [10], object detection for autonomous driving [21] etc. Since relying solely on visible images [40, 30] will render objects invisible under limited illumination, utilizing both visible (RGB) images and infrared (IR) images has been widely used as one of the effective solutions to achieve full-time object detection tasks. Specifically, RGB modality can provide the texture of objects in daylight, while infrared modality can provide the outline of objects under poor lighting conditions. By exploring the complementary information between the two modalities, RGB-IR object detection [15, 42] can aggregate the advantages of each modality to achieve robust visual perception.
@@
To design advanced fusion strategies for the MMOD task, previous works [17, 13, 29] reveal that the feature-level fusion methods outperform both image-level and decision-level fusion methods. Building upon this finding, recent RGB-IR object detection methods mainly focus on designing complex feature fusion structures to address various challenges. For example, CSSA [2] and ICAFusion [25] introduce different attention mechanisms to better explore multi-modal complementarity. CALNet [7] proposes the selected cross-modal fusion module to address the semantic conflict issue between different modalities. AR-CNN [43, 44] and TSFADet [35, 37] focus on solving the misalignment problem and attempt to improve object detection performance by aligning RGB and IR modality features. Although these methods have achieved encouraging object detection performance, they primarily focus on integrating complementary features from RGB-IR modalities while neglecting the mono-modality insufficient learning problem. This results in an unreasonable phenomenon that some objects can be detected by the mono-modal detector but missed by the corresponding multi-modal detector. We name this phenomenon as Fusion Degradation as illustrated in Figure 1. Besides, we statistics this phenomenon on three commonly used MMOD datasets and find that it is prevalent in different types of multi-modal detectors. Motivated by this, we rethink the multi-modal object detection task from the mono-modality learning perspective.
@@
In this paper, we construct M2D-LIF, an end-to-end framework for multi-modal object detection, to address the above problem. Different from the existing MMOD methods that design complex feature fusion modules, M2D-LIF facilitates the sufficient learning of mono-modality during multi-modal joint training and explore a lightweight yet effective feature fusion manner to achieve superior object detection performance. Inspired by knowledge distillation [8], we propose an Mono-Modality Distillation (M2D) method, which introduces the additional encoder pretrained on the mono-modality to distill the multi-modal encoders. To bridge the capability gap between the multi- and mono-modalities, we design the inner-modality and cross-modality distillation loss to jointly optimize the M2D-LIF during training. In this way, we can ensure sufficient learning of the mono-modal encoder in the multi-modality joint training. Besides, according to our observations in Section 3.1, we propose a Local Illumination-aware Fusion (LIF) module, which can dynamically set different weights for different illumination regions. The module ensures an explicitly complementary fusion to cooperate with M2D method, thereby enhancing both the accuracy and efficacy of the multi-modal object detection. Figure 1 shows that our framework can effectively solve the above problem. Our contributions in this paper are highlighted as follows:
@@
  • We introduce linear probing evaluation to the multi-modal detectors and identify the insufficient learning of mono-modality during training. To the best of our knowledge, it is the first time to rethink multi-modal object detection from a mono-modality learning perspective.
@@
  • We present M2D-LIF, a pioneering method to improve mono-modality learning capabilities during multi-modal joint training, innovatively exploring a lightweight yet effective feature fusion framework for the MMOD task.
@@
  • Extensive experiments on the three MMOD datasets demonstrate that our M2D-LIF outperforms the previous state-of-the-art detectors and can be used as an effective feature fusion method in the MMOD task.
@@ 2.1 RGB-Infrared Object Detection
However, the above methods mainly focus on designing complex and parameter-intensive modules to achieve the so-called complementary fusion, while ignoring the insufficient representation of each modality caused by the multi-modal learning. Different from these approaches, in this paper, we conduct the RGB-IR object detection task from mono-modality feature learning perspective and propose a novel M2D-LIF framework that only introduces a lightweight feature fusion module to achieve superior object detection performance.
@@ 2.2 Knowledge Distillation for Object Detection
Knowledge Distillation (KD) [8] can transfer knowledge from a high-capacity teacher model to a compact student model, thereby reducing the complexity of the model while ensuring model performance. For the object detection task, FRS [47] and PGD [32] leveraged foreground masks to enhance the ability of the student model to capture critical object features. CWD [26] normalized activation maps of each channel to obtain soft probability maps and minimized the KL divergence between these maps, enabling the student to focus on the most salient regions of each channel. Besides, PKD [1] imitated features with Pearson correlation coefficient to focus on the relational information from the teacher and relax constraints on the magnitude of the features. Recently, CrossKD [28] introduced cross-head knowledge distillation by transferring the intermediate features from the student detection head to the teacher providing a diversified perspective. As for the multi-modal object detection task, KD is usually used to transfer the knowledge from multi-modality features to mono-modality features. For example, distillation from Radar-Lidar to Radar [31], BEV-Lidar to BEV [6], and RGB-Infrared to RGB [20]. In our paper, unlike the above methods, we design a novel knowledge distillation method called Mono-Modality Distillation (M2D) to improve the feature representation of each backbone network in multi-modality joint training.
Figure 2: Liner probing evaluation on the FLIR dataset. Three types of feature fusion methods are selected for comparison, such as naive addition (Halfway Fusion [18]), weighted-based (IWM [9]), and attention-based (CMX [41]).
@@ 3.1 Linear Probing Evaluation
In the current MMOD framework, each modality is encoded by its corresponding backbone network, and then a fusion module is utilized to obtain the fused features for downstream perception tasks. Although achieving superior performance, we claim that this way will lead to insufficient learning of each modality and thus fail to achieve optimal detection performance. To validate this point, we employ the linear probing evaluation. Specifically, we first train the mono-modal object detectors and the multi-modal object detectors, respectively. For multi-modal object detectors, we select three popular feature fusion methods for comparison, such as Halfway Fusion [18] (naive addition), IWM [9] (weighted-based), and CMX [41] (attention-based). Note that, all detectors utilize the CSPDarknet53 [11] as the backbone network. After that, each backbone from both the mono- and multi-modality detectors is frozen and connected with the new detection head for training and testing. As shown in Figure 2, the evaluation results reveal that:
(1) The backbone networks from multi-modality joint training are insufficient learning. After linear probing evaluation, we can observe that all backbone networks from multi-modality joint training are worse than those from mono-modality training. This indicates that due to the existence of the fusion module, the learning ability of each backbone network is limited during multi-modal training.
@@
Figure 3: Overview of our $\rm M^{2}$-LIF framework. The proposed M2D is for enhancing mono-modal feature extraction capability of the backbone in the multi-modal object detection task. The proposed LIF method can evaluate the mono-modal quality based on RGB image’s local illumination conditions. Noted that since our framework is designed from the mono-modality perspective, the design of the infrared branch and the RGB branch is symmetrical and has the same network structure. Therefore, only the image processing flow of the RGB modality is emphasized in the figure, and the infrared branch is omitted.
@@ 3.2 M2D-LIF Framework
From Section 3.1, we have observed that the recent MMOD methods still suffer from the insufficient learning of mono-modality features. To solve this issue, referring to the knowledge distillation technology, we consider utilizing the additional encoder pretrained on the mono-modality as the teacher model to distill the multi-modality backbone network during training. Thus, under the premise that the multi-modal encoder learning sufficiently, we design a novel weighted-based feature fusion method to further improve the performance. The overall framework (M2D-LIF) is illustrated in Figure 3, which mainly consists of ➊ Mono-Modality Distillation and ➋ Local Illumination-aware Fusion. Note that the IR branch is omitted in Figure 3 since it has the same network structure as the RGB branch.
Mono-Modality Distillation (M2D). The M2D aims to enhance the feature extraction capability of the multi-modal encoders, laying a solid foundation for improving object detection performance. To bridge the capability gap between the multi- and mono-modal encoders, we first employ a pretrained teacher model of the same modality to distill the multi-modal backbone network using the inner-modality distillation loss $\mathcal{L}_{\text{IM}}$. To further enhance object-relevant feature extraction, we propose a cross-modality distillation loss $\mathcal{L}_{\text{CM}}$, which leverages cross-modality salient object location priors to guide feature distillation.
@@
The loss function $\mathcal{L}_{\text{IM}}$ represents the inner-modality feature distillation loss, which guides the backbone of the multi-modal object detection model to learn intermediate layer responses from a teacher backbone of the same modality. By minimizing $\mathcal{L}_{\text{IM}}$, the multi-modal backbone is encouraged to align with the feature responses of the teacher model. The loss is defined as follows:
@@
As for the loss function $\mathcal{L}_{\text{CM}}$, it represents the cross-modality feature distillation loss, which integrates cross-modality target location priors. Specifically, in MMOD tasks, different modalities generally contain information about the same objects. We first employ an attention mechanism to extract the salient object feature attention map, which serves as the location prior. The attention map is then used as a mask to guide object-relevant feature learning. To maintain a lightweight design, we adopt the parameter-free attention method SimAM [33]. The attention map $\widetilde{\mathcal{M}}$ can be calculated as:
@@
where $B$ denotes the predicted brightness map. We transform the RGB image to the $LAB$ color space and extract the $L$ channel as the ground-truth $\widetilde{B}$ to supervise the brightness prediction, which is formulated as follows:
@@
where $W_{V}$ and $W_{I}$ represent the weight of the RGB and infrared modalities, respectively. The hyperparameter $\alpha$ is the threshold that determines the importance of RGB features, and $\beta$ is the amplitude of the $W_{V}$. Thus, the boundary of $W_{V}$ and $W_{I}$ are:
@@
Finally, the LIF module uses the weight maps $W_{V}$ and $W_{I}$ to perform element-wise weighted fusion for the multi-modal features. Therefore, the final fused feature can be represented as:
@@
where $\lambda_{M^{2}D}$ and $\lambda_{LI}$ are the hyperparameters that control the balance between each loss.
@@
Table 1: Comparison of detection performance (mAP, in%) and computational cost (Params, FLOPs) between the LIF and other SOTA fusion methods on the FLIR (F), DroneVehicle (D), and LLVIP (L) datasets. The symbol indicates training with our M2D method. The Best results are highlighted by bold.
@@ 3.3 Why Using LIF Module for Feature Fusion
(1) robustness-and-reasonable. As analyzed in Section 3.1, the weighted-based method can serve as a competitive fusion way to improve performance. As a weighted-based fusion module, LIF is more robust to the insufficient learning of mono-modality encoders, thus achieving superior performance. Furthermore, unlike current weighted-based methods, the LIF module provides reasonably fine-grained quality assessment for the mono-modality images and explicitly achieves the complementary fusion.
(2) lightweight-yet-effective. As shown in Table 1, although attention-based methods can adaptively calculate the required features between modalities, they typically rely on high computational complexity. Unlike other methods, our LIF introduces only an additional 0.1M parameters and 6.4 GFLOPs in computational cost to achieve the superior performance, which is more lightweight and effective.